### v2

# LinkAlign: Scalable Schema Linking for Real-World Large-Scale Multi-Database Text-to-SQL

Abstract
Schema linking is a critical bottleneck in achieving human-level performance in Text-to-SQL tasks, particularly in real-world large-scale multi-database scenarios. Addressing schema linking faces two major challenges: (1) Database Retrieval: selecting the correct database from a large schema pool in multi-database settings, while filtering out irrelevant ones. (2) Schema Item Grounding: accurately identifying the relevant tables and columns from within a large and redundant schema for SQL generation. To address this, we introduce LinkAlign, a novel framework that can effectively adapt existing baselines to real-world environments by systematically addressing schema linking. Our framework comprises three key steps: multi-round semantic enhanced retrieval and irrelevant information isolation for Challenge 1, and schema extraction enhancement for Challenge 2. We evaluate our method performance of schema linking on the SPIDER and BIRD benchmarks, and the ability to adapt existing Text-to-SQL models to real-world environments on the SPIDER 2.0-lite benchmark. Experiments show that LinkAlign outperforms existing baselines in multi-database settings, demonstrating its effectiveness and robustness. On the other hand, our method ranks highest among models excluding those using long chain-of-thought reasoning LLMs. This work bridges the gap between current research and real-world scenarios, providing a practical solution for robust and scalable schema linking. The codes are available at https://github.com/Satissss/LinkAlign.
[…]
[backgroundcolor=mygray, linewidth=0pt ]newshaded
LinkAlign: Scalable Schema Linking for Real-World Large-Scale Multi-Database Text-to-SQL
Yihan Wang1thanks:    Authors contributed equally.  Peiyu Liu2* 1Beihang University  2University of International Business and Economics yihan3123@gmail.comliupeiyustu@163.com
1 Introduction
Figure 1: Comparision across methods on Spider 2.0-lite benchmark. Our method ranks highest, excluding those using long CoT reasoning LLMs, by simply applying LinkAlign to the DIN-SQL base model.
Text-to-SQL Zhong et al. (2017); Wang et al. (2017); Cai et al. (2017); Qin et al. (2022) aims to enable non-expert users to retrieve data effortlessly by automatically translating natural language questions into accurate SQL queries. Recent advances in large language models (LLMs) have led to notable improvements in Text-to-SQL benchmarks Sun et al. (2023); Pourreza et al. (2024), showcasing their growing capabilities in understanding and generating SQL queries. However, despite these advances, existing methods often struggle with real-world applications due to enterprise databases often containing massive redundant schemas and multi-database settings (e.g., local SQLite and cloud BigQuery). It faces significant failures in adapting existing Text-to-SQL models to large-scale multi-database scenarios, largely due to schema linking, i.e., identifying the necessary database schemas (tables and columns) from large volumes of database schemas for a user query Wang et al. (2019); Guo et al. (2019); Talaei et al. (2024). The underlying reasons for these failures remain unexplored, leaving a gap in addressing the real-world Text-to-SQL tasks.
To understand the failures, we perform a comprehensive error analysis and conclude that schema linking failures stem from two major challenges. Challenge 1 - Database Retrieval : how to select the correct database from a large schema pool in multi-database settings, while filtering out irrelevant ones. Existing researches often overlook this challenge as they always assume that single-database schemas are small-scale and can be directly fed into models for efficient processing. Challenge 2 - Schema Item Grounding : how to accurately identify the relevant tables and columns from within a large and redundant schema for SQL generation. The post-retrieval phase must handle complex database schemas, where the large volume of semantically similar tables and columns increases the risk of overlooking critical schema items necessary for generating accurate SQL queries.
Motivated by these factors, we propose LinkAlign, a novel framework that systematically addresses the challenges of schema linking in large-scale, multi-database scenarios. Specifically, our framework consists of three major steps. To address Challenge 1, our approach focuses first on (1) Retrieving Potential Database Schemas through multi-round semantic alignment by query rewriting. This step infers missing schema elements based on the retrieved result evaluation and rewrites the query to align with the missing database schema. Then our approach centers on (2) isolating irrelevant schema information to reduce noise by response filtering. This step filters the most appropriate database from a set of candidates, minimizing interference from irrelevant schemas, and streamlining the downstream processing pipeline. To address Challenge 2, our approach directs efforts towards (3) extracting schemas for SQL generation through identifying tables and columns by schema parsing. This step pushes the boundary of schema linking in large-scale databases by introducing advanced in-context learning techniques and carefully designed prompts for LLMs. To balance efficiency and accuracy, we introduce two distinct implementation strategies for our framework, i.e., pipeline mode and iterative mode. For the pipeline mode, a single LLM call is used to process each step, which is particularly well suited for scenarios where efficiency is critical. In contrast, the iterative mode adopts a multi-round multi-agent collaboration, which prioritizes accuracy and excels at handling complex and ambiguous queries, making it ideal for challenging Text-to-SQL tasks.
To better evaluate the model’s capabilities of schema linking, we construct AmbiDB, a variant of the Spider Yu et al. (2018) benchmark, designed to simulate the ambiguity and complexity of Text-to-Sql in real-world environment. We evaluate the schema linking capability of our approach on the SPIDER, BIRD Li et al. (2024a) benchmarks and the AmbiDB datasets proposed by ourselves, comparing it with existing baseline methods. Experimental results show that our approach achieves state-of-the-art performance across all metrics under multi-database settings. Additionally, we apply LinkAlign to the classic DIN-SQL method and conduct experiments on the SPIDER 2.0 Lei et al. (2024) benchmark to validate the effectiveness in adapting existing baselines to large-scale, multi-database scenarios. The results in Figure 1 show that LinkAlign ranks highest excluding those using long CoT reasoning LLMs.
2 Related Work
Early Work in Text-to-SQL tasks. Early approaches in Text-to-SQL systems primarily rely on rule-based methods Zelle and Mooney (1996); Saha et al. (2016), where manually predefined templates are designed to capture relationships between user queries and schema elements. The development of neural network-based approaches, particularly seq2seq architectures such as LSTM Hochreiter (1997); Sutskever (2014) and transformers Vaswani (2017) significantly improve text-to-sql performance Choi et al. (2021). Models like IRNet Guo et al. (2019) and RASAT Qi et al. (2022) leverage attention mechanisms to better incorporate schema elements into the query understanding process. A notable breakthrough in schema linking is achieved with the introduction of graph neural networks Wang et al. (2019); Cao et al. (2021); Bogin et al. (2019), enabling models to represent relational database structure as graphs and learn connections between query and schema elements more effectively.
Evolution Techniques elicited by LMs. With the development of pre-trained language models such as BERT Devlin (2018) and T5 Raffel et al. (2020) , approaches such as SADGA Cai et al. (2021) and PICARD Scholak et al. (2021) combine PLMs with task-specific fine-tuning to enhance execution accuracy. Recent studies also explore the integration of LLMs Rajkumar et al. (2022); Liu et al. (2023); Guo et al. (2023), such as GPT, which can directly generate SQLs without requiring task-specific training data. Building on LLMs, models like C3 Dong et al. (2023), DIN-SQL Pourreza and Rafiei (2024), and MAC-SQL Wang et al. (2024) leverage task decomposition strategies and advanced reasoning techniques, such as Chain-of-Thought Wei et al. (2022), Least to Most Zhou et al. (2022) and self-consistency decoding Wang et al. (2022) to address schema linking tasks more effectively.
Emerging Solutions and Challenges. Schema linking remains a challenge when handling large-scale schema and further complicated by the ambiguity in user queries. Approaches such as CHESS Talaei et al. (2024) and MCS-SQL Lee et al. (2024) try to address this by using multiple intricate prompts and sampling responses from LLMs on existing benchmarks such as Spider Yu et al. (2018) and Bird Li et al. (2024a). Although effective, they still require considerable computational resources and API costs, particularly in large-scale database scenarios.
3 Error Analysis
To better understand the gap between existing research and real-world environment, we test 500 samples from the Spider dataset and analyze common error types in existing baseline methods for Text-to-SQL tasks. In particular, the model needs to handle schemas from all databases, rather than small-scale schemas from a single database, which is more aligned with real-world scenarios. The results show that schema linking errors are the main cause of Text-to-SQL failures, with an error rate greater than 60%. To better understand the failures, we manually examined the erroneous samples and attributed the four common error types to two major challenges proposed in the Introduction. Appendix A provides more details of the results.
Error 1: Target Database without Retrieval (Database Retrieval). This error indicates that the retrieved results do not include complete ground-truth database schemas, accounting for 23.6% of the failures. For example, the user intends to query "which semester the master and the bachelor both got enrolled in", but the retrieved results exclude $degree\_program$ table from the target database, which requires inference based on query semantics and commonsense knowledge. However, general vectorized retrieval approaches only return semantically similar results based on embedding distance, conflicting with the fact that user complex queries often misaligned with the ground-truth schema.
Error 2: Referring Irrelevant Databases (Database Retrieval). Unlike Error 1 that focuses on the missing target database, Error 2 centers on the irrelevant database noise introduced by imprecise retrieval under multi-database settings. This error indicates that the model refers to incorrect schemas from unrelated databases when generating SQL, accounting for 13.3% of the failures. For example, the user intends to query "The first name of students who have both cats and dogs". However,the generated SQL incorrectly infers $People.first\_name$ from an unrelated database while overlooking the ground truth $Student.f\_name$, even though both are successfully retrieved and fed into the model. Without isolating irrelevant database information, the model tends to mistakenly select seemingly more appropriate schemas from unrelated databases, leading to SQL execution failures.
In summary, Error 1 and Error 2 highlight the gap between existing methods and real-world large-scale multi-database scenarios. Even though these are avoided, SQL execution still fails due to other schema linking errors.
Error 3: Linking to the Wrong Tables (Schema Item Grounding). This error indicates that the generated SQL omits or misuses tables, accounting for 19.8% of the failures. For example, both the $student$ and $people$ tables have fields for $name$, but the model selects the latter incorrectly. In real-world scenarios, the models often overlook critical tables, which directly impairs the execution accuracy of the generated SQL Maamari et al. (2024).
Error 4: Linking to the Wrong Columns (Schema Item Grounding). This error indicates that the generated SQL omits or misuses fields despite referencing the ground truth table correctly, accounting for 11.6% of the failures. For example, the model may omit the join columns $pets.pet\_id$ and $has\_pet.pet\_id$ in the join operation of the correct SQL statement. Missing such critical columns directly incurs SQL execution failure.
Figure 2: Overview of the LinkAlign framework including three core components.
4 Methodology
This section presents the LinkAlign framework for schema linking and the implementation strategies used to realize it. Our framework consists of three major steps designed to address the challenges in large-scale, multi-database scenarios. The implementation strategies provide two complementary approaches to execute these steps balancing efficiency and accuracy.
4.1 Background
Before proposing our method, we consider a typical Text-to-SQL setting. Given a set of $N$ databases $D=\{D_{1},D_{2},\dots,D_{N}\}$ and the schemas $S=\{S_{1},S_{2},\dots,S_{N}\}$, where a schema $S_{i}$ is defined as $S_{i}=\{T_{i},C_{i}\}$, with $T_{i}$ representing multiple tables $\{T_{1}^{i},T_{2}^{i},\dots,T_{|T_{i}|}^{i}\}$ and $C_{i}$ representing columns $\{C_{1}^{i},C_{2}^{i},\dots,C_{|C_{i}|}^{i}\}$. Traditional methods take full multi-database schemas and user query as input and rely on schema linking component to identify tables and columns for SQL generation:
$S^{\prime}=f_{\text{parser}}\left(S,Q,c\mid E,LLM\right),$ (1)
where $f_{\text{parser}}\left(\cdot\mid E\right)$ denotes the schema parsing function based on the text-embedding model $E$ and LLM. Symbol $c$ denotes additional context such as field descriptions or sampling examples.
4.2 LinkAlign Framework
The LinkAlign framework can effectively adapt existing models to real-world environments by systematically addressing the challenges of schema linking. The framework consists of three main steps that facilitate the downscaling of database schemas and enhance schema linking capabilities.
Step one: retrieve potential database schemas. To mitigate the Error 1, we propose a method for multi-round semantically enhanced retrieval. In this step, LLMs are used to evaluate the retrieval results and rewrite the input query based on the inferred missing database schemas. Dynamically adjusting the retrieval strategy to align semantically with missing database schemas enables high retrieval performance with fewer iterations. This approach is particularly effective in scenarios with limited retrieval batches and restricted scale. {newshaded}User Query $Q_{0}$: Which semester the master and the bachelor both got enrolled in?
Missing Schemadegree_programs (degree_type) [1] Rewrite $Q_{1}$: In a database with degree_programs, how to find semesters where both master’s degree_type and bachelor’s degree_type programs exist? Group by enrollment_semester with checks for both program types.   Missing Schemaenrollment_records (semester)
[2] Rewrite $Q_{2}$: In a database with enrollment_records, how to find semesters where both master and bachelor students enrolled? Group by semester and filter for overlapping enrollments. The rewritten queries are embedded into vectors and then used to retrieve the relevant database schemas. The retrieval results are sorted and summarized based on the number of rewrites and similarity scores:
$Z=\bigcup_{t=0}^{T}f_{\text{retriever}}\left(S,Q_{t},c\mid E\right),$ (2)
where $T$ represents the number of query rewrites. By rewriting queries and enhancing their semantic representation, this approach improves the alignment between user queries and database schemas, ensuring more accurate retrieval outcomes.
Step two: isolate irrelevant schema information. To mitigate Error 2, we attempt to filter out irrelevant schemas from the retrieved multi-database results. Once the retrieved results $Z$ containing schemas from irrelevant databases, the next step is to precisely locate the target database $D_{t}$ while filtering out irrelevant ones. This component is designed to minimize interference from extraneous schemas, which can obscure the generation of accurate SQL queries. To achieve this, the framework evaluates the relevance of each database $D_{i}$ from retrieved results $Z$ based on the original user query $Q_{0}$. The most likely target database $D_{t}$ is selected based on its analyzed alignment with the query.
$D_{t}=\arg\max_{1<i<N}P_{M}(D_{i}\mid Q_{0},Z),$ (3)
where M denotes the LLMs used for analysis. By focusing on the most appropriate database, this step isolates unrelated schemas information from being included in subsequent operations. This filtering is essential, as irrelevant schemas can unnecessarily extend the input context, increasing computational costs and diminishing SQL generation performance. By selecting only the target database, the framework ensures that schema linking focuses on relevant information, thereby improving both accuracy and efficiency. Additionally, our approach remains effective in single-database settings by adjusting the objective to filter out irrelevant database schemas rather than those from unrelated databases, as explained in Appendix C.
Step three: extract schemas for SQL generation. To mitigate the Error 3 and 4, we need to identify the correct tables and columns for SQL generation. This step mimics the manual process of database schema extraction, but it can be automated using reasoning techniques such as Chain-of-thought. Specifically, given filtered database schema $S_{\hat{u}}$ from previous steps, we need to identify a subset $S_{\hat{u}}^{\prime}$ that contains the most relevant tables $T_{i}^{\hat{u}}$ and columns $C_{i}^{\hat{u}}$ These elements are selected based on their alignment with the query, ensuring that the resulting schema is both precise and comprehensive.
$S_{\hat{u}}^{\prime}=\{T_{i}^{\widehat{u}},C_{i}^{\widehat{u}}|\ \mathbb{I}(Q_{0},C_{i}^{\widehat{u}})=1\}\},$ (4)
[…]
4.3 Component Implementation Optimization
Based on the formatted definitions of each step in Section 4.2, we propose two strategies to further implement the key components in each step, with positions shown in the dashed box in Figure 2. These strategies share modular interfaces and the specific calls tailored to different levels of query complexity and database scale. The first one is Single-Prompt Pipeline, which simplifies the implementation of the LinkAlign framework through a single call to LLMs. This approach is particularly well-suited for large-scale scenarios where efficiency is critical. Appendix B provides a detailed explanation of this strategies, while this section focuses on Multi-agent Collaboration strategy, which prioritizes accuracy and provide sufficient capabilities to tackle complex query tasks in real-world environment.
Align Semantics by Query Rewriting. Inspired by the reflection capability of LLMs discussed by Shinn et al., we design an semantic enhanced retrieval approach composed of two step: inferring missing schemas and query rewriting. Specifically, the Schema Auditor first evaluates the retrieval results, infers the missing schemas required for accurate SQL generation (e.g., fields for SELECT/JOIN/WHERE), and concludes with an audit report. Next, the Query Rewriter enriches the question with additional semantic information to better align with ifferred missing database schemas, which ultimately enhances retrieval accuracy.
Reduce Noise by Response Filtering. When multiple candidate databases show minor differences, reaching consensus through multi-agent debates significantly reduces the risk of errors Chan et al. (2023). Inspired by this, we develop a multi-agent debate model composed of Data Analyst and Database Expert. Specifically, the Data Analyst infers the most appropriate database from multiple candidates, and the Database Expert evaluates whether the selection is reasonable and decides retain or not. The debate follows a one-by-one strategy, i.e., starting with the data analyst, after which the two roles present their perspectives in turn. The debate ends when a predefined number of rounds is reached, and then a terminator outputs the consensus database as a final result.
Identify Tables and Columns by Schema Parsing. To enhance schema linking capabilities, we develop a multi-agent debate model based on Simultaneous-Talk-with-Summarizer strategy, i.e., multiple peer participants engage in simultaneous debate, with the final outcomes evaluated by a summarizer. The debate involves two roles: Schema Parser and Data Scientist. In each round, each Schema Parser extracts a clause-specific database schema (e.g. fields for JOIN/WHERE…), and then the Data Scientist verifies all results from Schema Parsers, identifying any omissions or errors. Multi-role participation enhances the recall of tables and columns required for SQL generation, with diverse answers complementing each other to reduce the randomness of single-prompt outputs.
Approach Spider Bird AmbiDB
LA EM Recall LA EM Recall LA EM Recall
LlamaIndex
DIN-SQL 80.0 26.8 62.4 68.8 5.1 31.3 59.7 13.3 44.2
PET-SQL 84.1 38.6 67.2 77.1 8.2 39.7 66.4 22.0 50.2
MAC-SQL 82.3 17.3 42.8 75.0 5.7 34.5 65.1 9.7 30.8
MCS-SQL 81.0 24.3 73.2 73.7 13.9 56.1 61.9 13.7 54.8
RSL-SQL 74.8 29.1 76.1 80.0 16.1 61.8 62.4 17.9 59.6
Pipeline(ours) 85.4 37.4 65.9 66.8 8.6 38.1 69.4 20.3 50.4
Agent(ours) 86.4 47.7 80.7 83.4 22.1 64.9 67.6 22.4 56.9
Table 1: Comparison of LA, EM and Recall across different methods in multi-database scenario.
Approach Spider-dev Bird-dev AmbiDB
Precision Recall EM Precision Recall EM Precision Recall EM
DIN-SQL 83.9 73.2 40.4 79.9 55.7 13.1 86.6 76.9 44.2
PET-SQL 84.8 73.9 33.4 81.6 64.9 25.9 90.2 78.3 39.5
MAC-SQL 75.0 66.8 24.4 76.3 56.2 9.1 79.8 69.6 30.1
MCS-SQL 66.7 85.0 29.8 79.6 76.9 25.5 71.5 88.5 34.1
RSL-SQL 74.8 84.3 37.6 78.1 77.5 27.7 80.7 88.3 42.2
Agent (ours) 80.2 87.3 48.1 77.1 79.4 29.0 86.7 85.8 51.5
Table 2: Comparison of Precision, Recall and EM across different methods in single-database scenario.
5 Experiments
5.1 Experiemental Setup
Dataset. We evaluate our method performance of schema linking on the SPIDER, BIRD and AmbiDB datasets, and the ability to adapt existing Text-to-SQL models to real-world environments on the SPIDER 2.0 benchmark. We provide more details about SPIDER, BIRD and SPIDER 2.0 in Appendix  E, and the construction of the AmbiDB dataset in Appendix  D.
Baselines.We compare our method with the following baselines.
$\bullet$ DIN-SQL Pourreza and Rafiei (2024) introduces a prompt-driven schema linking method with a single LLM call which integrates a chain-of-thought strategy to enhance reasoning.
$\bullet$ PET-SQL Li et al. (2024b) leverages LLMs’ SQL generation capabilities to generate preliminary SQL (PreSQL) statements for deriving schema references.
$\bullet$ MAC-SQL Wang et al. (2024) employs a Selector agent to identify the minimal relevant schema subset with a single LLM call.
$\bullet$ MCS-SQL Lee et al. (2024) employs a two-step process for table and column linking, utilizing multiple prompts and random shuffling to enhance recall and robustness.
$\bullet$ RSL-SQL Cao et al. (2024) introduces a bidirectional schema linking approach that integrates both forward and backward linking to streamline the database schema.
Metrics. We evaluate the ability of schema linking using the following metrics:
$\bullet$ Locate Accuracy (LA). Let $N$ denote the total number of test examples and $N_{a}$ the number of examples without Error 1 or Error 2. The LA is defined as $N_{a}$ / $N$, measuring the model’s ability to locate the target database accurately.
$\bullet$ Exact Matching (EM). Let $N_{e}$ denote the number of examples without Error 1 to 4. The EM score is defined as $N_{e}$ / $N$, measuring the model’s ability to perform precise schema linking.
$\bullet$ Recall. This metric measures the recall of database schemas from the ground truth SQL. It is preferred over Precision, as minor schema noise may not significantly impact SQL generation Maamari et al. (2024). But it still makes sense for the model to maintain a high recall rate while improving precision, in order to minimizing noise introduced by excessive irrelevant schema.
We evaluate the ability of Text-to-SQL using the Execution Accuracy:
$\bullet$ Execution Accuracy(EX). This metric is widely used to evaluate the quality of the generated SQL Yu et al. (2018); Li et al. (2024a); Lei et al. (2024), based on the comparison with the results of the Gold SQL execution.
Hyperparameter. The open-sourced text-embedding model bge-large-en-v1.5 is applied to convert database schema metadata and queries into vectors. We set the top-K of retrieval size at 5 and we use GLM-4-air model for experiments.
5.2 Main Results
Results on Multi-Databases. As shown in Table  1, our method achieves state-of-the-art Locate Accuracy (LA) on the SPIDER, BIRD, and AmbiDB datasets, with scores of 86.4%, 83.4%, and 69.4%, respectively, demonstrating the framework’s effectiveness in mapping data requirements to the target database. Unlike baseline methods, our framework introduces the "Reduce Noise by Response Filtering" step that filters out irrelevant database schemas, reducing error 2 and maintaining high LA accuracy. This step also benefits the model by eliminating the need to choose between similar candidate schemas, allowing it to identify the required tables and columns more accurately. Experimental results confirm this: our framework achieves state-of-the-art Exact Match (EM) performance across all three datasets, surpassing the baseline models by 23.6%, 1.8%, and 37.3%, respectively. The presence of error 1,2 in multi-database contexts makes it difficult for the model to avoid blending unrelated schemas, leading to challenges in accurately recalling relevant tables and columns. Especially as the database size increased, we observed a significant decrease in the recall rate of all methods on the AmbiDB dataset, further demonstrating that our proposed dataset exacerbated the challenge.
Results on Single Database. We further compare and evaluate the ability of different methods to identify correct tables and columns in a large-scale database. As shown in Table  2, our Agent method achieves state-of-the-art performance across all datasets in terms of the EM evaluation metric. After eliminating interference from irrelevant database schemas, the recall rate for all methods significantly improved compared to the results in Table  1. Compared to the baseline models, the Agent method achieves the highest recall rates on both the Spider-dev and Bird-dev datasets. This demonstrates that our method exhibits superior performance and robustness when the inference capabilities of large models are limited. Although, on the AmbiDB dataset, our method’s recall rate lags behind MCS-SQL (88.5%) and RSL-SQL (88.3%), our method outperforms these models by 21.3% and 7.4%, respectively, in Precision. This indicates that our method maintains high recall while minimizing irrelevant noise. Overall, considering all metrics, our method demonstrates excellent performance and robust schema-linking capabilities.
Model variant Spider AmbiDB
LA EM Recall LA EM Recall
Pipeline 85.4 37.5 66.1 69.4 20.3 50.4
w/o que. rew. 85.3 37.7 72.3 63.1 14.5 52.8
w/o res. fil. 81.9 26.0 62.0 66.2 15.3 48.5
w/o both 80.0 26.8 62.4 59.5 11.4 38.7
Agent 86.4 47.7 80.7 67.6 22.4 56.9
w/o que. rew. 83.6 30.6 73.0 65.3 15.1 57.0
w/o res. fil. 66.7 27.8 54.8 58.5 14.5 60.6
w/o both 73.6 32.9 61.1 58.0 17.6 47.8
Table 3: Performance comparison of model variants on Spider and AmbiDB datasets. “que. rew.” indicates query rewriting and “res. fil.” denotes the response filtering.
Figure 3: The impact on Error Rates.
Figure 4: The impact on Evaluation Metrics.
Results on Advanced Spider 2.0-lite Benchmark. To convincingly demonstrate the effectiveness of our method, we conducted tests on Spider 2.0 Lei et al. (2024), an open-source benchmark which is tailored for evaluation on real-world challenges. Our submission results are shown in the Figure 1. Notably, our method using Qwen-PLUS significantly outperforms other models like DailSQL, CHESS, and DIN-SQL using GPT-4o, highlighting the effectiveness of our approach. Furthermore, our approach, LinkAlign+DeepSeek-v3 ranks overall 4th place across all sequences without introducing test-time scaling, in contrast to the top three methods which all employ OpenAI O1 and O3 models. This underscores the potential of our method for further improvement in the future.
5.3 Ablation Study
We conducted a ablation study to investigate incremental impact of two core step of LinkAlign framework.The results are presented in Tables  3 , demonstrated that each step of proposed methods contributes to the SOTA performance on benchmark datasets.
Impact of Query Rewriting. User queries often fail to align with their required schema, leading to inefficient retrieval. As shown in Figure  4, introducing Query Rewriting step reduced error 1 rates of two strategies by 6.9% and 10.8%, respectively. This demonstrates this step effectively eliminates ambiguity and improves retrieval recall of target database.Compared to Pipeline , the Agent strategy shows more significant benefits, suggesting LLMs reflextion more effectively aligns the query with the schema. However, this step also introduces more irrelevant database schemas, which increases the challenge of database-locate and leads to a rise in error 2 by 5.7% and 8.4%, respectively.Since the slope of the negative impact from error 2 is lower than the positive gain from solving error 1, the Locate Accuracy metric shows an overall improvement. Compared to the Spider benchmark, the LA improvement on the AmbiDB dataset is more pronounced, indicating this step plays a more critical role when ambiguity in queries increases. For simple query scenarios, it may be necessary to balance the positive gains and negative impacts of query rewriting to maximize the LA benefits.
Impact of Response Filtering. As shown in Figure  4, although Query Rewriting introduces many irrelevant databases, the Response Filtering step reduces the error 2 rate of the Agent strategy by 10.8%, effectively mitigating the negative impact of the previous step. Compared to the Pipeline strategy, the Agent strategy benefits more significantly, indicating that this step plays a more critical role within the Agent strategy. Filtering allows LLMs to focus on single-target database schema more narrowly, making it easier to accurately identify the most relevant schema. As shown in Table  3, Response Filtering effectively improves both the EM and Recall metrics for both strategies. Although selecting the wrong database directly causes failure in Schema Linking task, Response Filtering significantly improves schema linking performance by mitigating the negative impact of error 3,4.
6 Conclusion
In this paper, we aim to adapt existing methods to real-world large-scale multi-database scenario by tackling the challenge of schema linking. First, we highlight four core errors leading to schema linking failures. Based on this analysis, we propose the LinkAlign framework, which composes of three key steps. Additionally, we introduce the AmbiDB dataset, for better design and evaluation of the schema linking component. Experiments demonstrate that our model outperforms existing baseline methods across all evaluation metrics in both multi-database and single-database contexts.
7 Limitations
This paper presents two limitations that will require attention in future work. First, we did not explore the potential advantages of combining different strategies for addressing the schema linking task, even though our modular interface could facilitate such combinations. Second, we did not use the most advanced LLMs in our experiments. The reasoning limitations of the LLMs maybe resulte in more noticeable performance between different methods. However,as large models gain sufficient capabilities, we may need to consider whether it’s necessary to simplify the framework.
Appendix A Error Analysis
To figure out why existing methods fail in real-world environments, we tested 500 examples randomly sampled from the SPIDER dataset. In particular, the model needs to handle schemas from all databases, rather than small-scale schemas from a single database. Given the challenges existing methods face in handling large-scale database schemas, we adopt a vectorized retrieval approach to search relevant schemas based on user queries. Then the retrieve results composed of related schemas from different databases are fed into DIN-SQL models to generate SQL for user queries.
Experimental results in Figure 5 show that in large-scale, multi-database scenarios, the EX score of the DIN-SQL model drops from 85.3 to 67.4%. Manual analysis of error cases reveals that schema linking errors account for 68.3% of Text-to-SQL failures, making them the major failure cause.
Figure 5: Error Distribution in Failed Cases.
Appendix B Single-Prompt Pipeline Strategy
This section provides a detailed introduction of the single-prompt pipeline strategy which simplifies the LinkAlign framework for better efficiency.
Align Semantics by Query Rewriting.We propose a query semantic enhancement module that utilizes few-shot Chain of Thought examples to guide the LLMs to clarify the query’s semantic intent through four reasoning steps.
Step 1: Requirement Understanding. The first step involves rephrasing the user query to explicitly define its objective and data requirements.
Step 2: Key Entity Identification. This step extracts and identifies the key entities or values from the query that are semantically relevant to the target database.
Step 3: Entity Classification. Based on the previous step’s extractions, entities are classified into broader categories, and their relationships are defined.
Step 4: Database Schema Inference. The final step infers the relevant tables and columns in the target database schema that are likely to provide the necessary data to address the query.
Reduce Noise by Response Filtering. We design a prompt with few-shot Chain of Thought examples to guide the LLM through multiple reasoning steps, mapping the query to the correct target database. First, the model rephrases the data requirements to ensure full understanding. Next, it evaluates each database to confirm it contains the necessary data columns. Finally, the model outputs the name of the most relevant database.
Identify Tables and Columns by Schema Parsing. We adapted the prompt design from DIN-SQL, employing a single LLM call for schema linking. DIN-SQL uses a prompt with few-shot Chain of Thought examples, incorporating 10 randomly selected samples from the Spider dataset’s training set. Experimental results demonstrate that this approach strikes a balance between accuracy and efficiency in database query scenarios with a small number of tables and columns, aligning with the objectives of the pipeline method. By utilizing a single model call, the method achieves strong performance in solving simpler problems.
Appendix C Effectiveness in Single-Datbase Setting
In this section, we discuss how to effectively apply the LinkAlign framework to Text-to-SQL models in large-scale single-database scenarios. As a common assumption in existing studies, the single database setting simplifies real-world multi-database environments by only utilizing the schema from target database as input for Text-to-SQL models. However, the single-database setting is often impractical in real-world scenarios, as non-technical business users always struggle to select the appropriate database for data query (e.g., local SQLite and cloud BigQuery) due to a lack of expertise in database architecture. This is why our study prioritizes the multi-database setting, as this gap limits the real-world application of existing models.
To ensure LinkAlign remains effective in the single-database setting, we need to adjust the objective of Step two to filter out irrelevant database schemas rather than those from unrelated databases. However, we notice that this approach sometimes mistakenly excludes the ground-truth schema, as it may appear unrelated to generating the correct SQL. We propose two feasible optimization techniques to address this issue as explained below. Specific adjustments of prompts and codes are available in our open source repository.
Random preservation with exponential decay. To avoid discarding potentially correct schemas without sufficient evidence, we randomly retain database schemas for each retrieval round using a dynamic retention rate. Specifically, the retention rate decays exponentially with the retrieval rounds because rewritten queries may gradually deviate from the user’s original intent, thereby increasing noise. Random sampling of retrieval results not only preserves expected benefits, but also enhances the method’s generalization ability. In addition, the retention rate needs carefully chosen to ensure that the number of retained schemas is smaller than the excluded ones. In our experiments, we set the initial retention rate (at turn n = 0) to 0.55, the exponential decay coefficient to 0.6, and clip it to 0 when it falls below 1.
Post Retrieval. This method is also highly effective by performing mini-batch retrieval on excluded database schemas (i.e., those not retrieved or filtered out). Intuitively, this offers the model a new opportunity to sift gold from the sand without the influence of the spotlight, as it compares against noisy database schemas rather than those obviously relevant. This stage employs the same method as Step One, differing only in the mini-batch retrieval scale and the number of retrieval rounds. In our experiments, we set the post-retrieval top-K to 5 and turn-n to 1.
Appendix D AmbiDB Dataset Construction
We introduce the AmbiDB benchmark, a variant of the Spider dataset. It is constructed through database expansion and query modification to better simulate real-world query scenarios characterized by large-scale synonymous databases and enhanced-ambiguity queries by multi-database contexts. The motivation stems from three limitations in existing benchmarks. First, experiments on the existing benchmarks specifies the target database required by the query in advance, ignoring the challenge of mapping user queries to the target database in multi-database scenarios. Second, the existing benchmark has a limited number of synonym databases, exhibiting lesser ambiguity in multiple databases context. Third, existing benchmarks often fail to balance database scale and variety. AmbiDB outperforms Spider in terms of scale and surpasses Bird in terms of quantity.
D.1 Data Construction
Database Expansion We extend the database schema through two key steps. First, we instruct the LLMs to extract a subset of schema from original database.The extracted schema subset forms the foundation to construct synonym databases. These schema subsets typically capture key characteristics of the specific domain. For example, the Student table, containing attributes such as student ID and name, can serve as a shared schema for synonym databases related to College domain. Second, we add new tables and columns to expand the scale of every database while aligned with the original database themes. The expanded schemas preserve the integrity of the original SQL queries while becoming larger in scale. Furthermore, the inclusion of similar sub-schemas across multiple databases enhances contextual ambiguity, making it more close to real-world query scenarios.
Query Modification We instruct the LLMs to modify original queries with the use of synonymous databases, which making them more complex and ambiguous in multi-database scenarios. The motivation for this step is that the original queries are semantically explicit or contain sufficient details, facilitating identifying required tables and columns from the target database. This step aligns the queries semantically with overlapping schemas included in multiple synonymous databases, while avoiding tokens with identical names across databases. Additionally, we subtly incorporated details that could be reasoned to help determine the target database, preventing complete confusion.
D.2 Data Filter
In the main text, we provided a brief overview of Query Modification. However, the complete process involves two key steps: (1) generating the correct SQL query for a given question, and (2) filtering the rewritten question-SQL pairs.
Generating the Correct SQL Query for the Question. LLMs with strong reasoning and generation capabilities is employed to generate the corresponding SQL query based on the modified question. Due to the inherent ambiguity and complexity of the questions, the model cannot guarantee absolute correctness of the generated SQL. To address this, we implement an automatic verification step, where the model checks if the data requirements described in the question align with the SQL query’s execution result. The model is then given a single opportunity to correct any errors by fine-tuning the question to match the SQL query. Finally, we manually review and verify the question-SQL pairs to ensure correctness.
Filtering the Rewritten Question-SQL Pairs. We assess the complexity and ambiguity of each question individually, removing or modifying any inadequate samples. We then filter out the top 10% and bottom 5% of questions based on length. This approach is driven by two considerations: first, longer questions may contain more semantically relevant information about the target database, making it easier to locate the question through semantic matching alone, which fails to simulate the real-world challenges of database localization. Second, shorter questions may oversimplify the Schema Linking task, making it easier to address the second challenge.
Appendix E Supplementary Experimental Setup
SPIDER includes 10,181 questions and 5,693 SQL queries spanning 200 databases, encompassing 138 domains. The dataset is split into 8,659 training examples, 1,034 development examples, and 2,147 test examples. The databases utilized in the multi-database scenario experiment are primarily sourced from the dev dataset and the train dataset.
BIRD comprises 12,751 question-SQL pairs across 95 large databases and spans 37 professional domains. It adds external knowledge to align the query with specific database schemas. The queries in BIRD are more complex than SPIDER, necessitating the provision of sample rows and external knowledge hints to facilitate schema linking.
SPIDER 2.0 comprises 632 real-world text-to-SQL tasks from enterprise databases with thousands of columns and diverse SQL dialects. It requires schema linking, external documentation, and multi-step SQL generation, posing greater challenges than SPIDER 1.0 and BIRD. Simplified variants, Spider 2.0-lite and Spider 2.0-snow, focus on text-to-SQL parsing without workflow interaction.
Appendix F SPIDER 2.0-lite Evaluation
We evaluate our method on the SPIDER 2.0-lite Benchmark. In the experiments, we applied LinkAlign to the DIN-SQL base model to adapt it to real-world large-scale query scenarios. Our method rank highest excluding those using long CoT reasoning LLMs. The results demonstrate that our method has a significant impact.
Figure 6: Performance of LinkAlign framework on SPIDER2.0-lite benchmark.